OpenStudy (anonymous):

The sum of two numbers is 48. If the larger number is 6 less than twice the smaller number, find the two numbers. Write and solve a system of equations for this situation.

OpenStudy (epoweritheta):

let the larger number =x ; y=smaller number x=2y-6 ; x+y=48; solving ; 48-y=2y-6 3y=54 y=18

OpenStudy (anonymous):

thanks!